# Neurons Speak in Ranges: Breaking Free from Discrete Neuronal Attribution

### v1

@@ Abstract
Interpreting and controlling the internal mechanisms of large language models (LLMs) is crucial for improving their trustworthiness and utility. Recent efforts have primarily focused on identifying and manipulating neurons by establishing discrete mappings between neurons and semantic concepts. However, such mappings struggle to handle the inherent polysemanticity in LLMs, where individual neurons encode multiple, distinct concepts. This makes precise control challenging and complicates downstream interventions. Through an in-depth analysis of both encoder and decoder-based LLMs across multiple text classification datasets, we uncover that while individual neurons encode multiple concepts, their activation magnitudes vary across concepts in distinct, Gaussian-like patterns. Building on this insight, we introduce $\mathsf{NeuronLens}$, a novel range-based interpretation and manipulation framework that provides a finer view of neuron activation distributions to localize concept attribution within a neuron. Extensive empirical evaluations demonstrate that $\mathsf{NeuronLens}$ significantly reduces unintended interference, while maintaining precise control for manipulation of targeted concepts, outperforming existing methods.
Polysemanticity,Interpretability
@@ 1 Introduction
Large language models (LLMs) have shown unprecedented performance in various tasks involving natural language understanding, generation, and transformation (Brown et al., 2020; Bommasani et al., 2022; Touvron et al., 2023; Raffel et al., 2019). However, the inner workings of LLMs remain largely opaque (Burkart & Huber, 2021), as their representations are distributed across billions of parameters. This lack of interpretability raises critical concerns about reliability, fairness, and trustworthiness, particularly in high-stakes domains such as healthcare, law, and education. To this end, neuron-level interpretability can address these concerns by enabling researchers to uncover how individual neurons encode semantic concepts and contribute to model outputs. With this understanding, researchers can diagnose safety risks (Wei et al., 2024; He et al., 2024), refine model outputs (Meng et al., 2023; Rizwan et al., 2024), optimize efficiency through pruning (Frankle & Carbin, 2019; Haider & Taj, 2021), and steer models toward desired objectives (Subramani et al., 2022).
Recent research efforts have made significant progress in neuron-level interpretation by identifying salient neurons that influence model behavior (Dalvi et al., 2019a; Antverg & Belinkov, 2022; Conmy et al., 2023; Marks et al., 2024). Approaches such as maximal activation analysis identify neurons that exhibit the highest activation values for specific concepts (Foote et al., 2023; Frankle & Carbin, 2019). Probe-based methods employ auxiliary classifiers to distinguish between concepts using hidden representations (Dalvi et al., 2019a, b), while probeless approaches bypass the need for classifiers by directly analyzing neurons, improving simplicity and interpretability (Antverg & Belinkov, 2022). Other techniques include circuit discovery, which attributes concepts to groups of interacting neurons (Olah et al., 2020; Conmy et al., 2023), and causal analysis, which intervenes on internal components to identify their role in model behavior (Vig et al., 2020).
Figure 1: $\mathsf{NeuronLens}$ leverages distinct, Gaussian-like activation patterns of neurons to enable fine-grained concept attribution and selective concept manipulation with minimal interference with other concepts.
While these methods have advanced our understanding, they often rely on discrete neuron-to-concept mappings, which assume that entire neurons encode single concepts. However, neurons frequently exhibit polysemanticity – the ability to encode multiple, seemingly unrelated concepts (Lecomte et al., 2024; Marshall & Kirchner, 2024). Given this heterogeneous encoding of concepts, traditional approaches often lead to unintended consequences when manipulating neurons, as changes intended for one concept may inadvertently affect others encoded by the same neuron or suboptimal interpretations of concepts (Sajjad et al., 2022).
@@
This work seeks to address these questions by analyzing the activation patterns of neurons in both encoder-based and decoder-based LLMs. Through statistical and qualitative analysis, we find that neuronal activations for concepts follow Gaussian-like distributions, with distinct patterns for different concepts. Our key insight is that the true fundamental unit of interpretability lies at a level more fine-grained than the neuron itself. Within a neuron’s activation spectrum, activation ranges corresponding to specific concepts emerge as the fundamental unit. This novel perspective enables a more precise approach to neuron interpretation and manipulation, addressing the limitations of traditional, discrete neuron-to-concept mappings.
Building upon these insights, we introduce $\mathsf{NeuronLens}$, a novel range-based framework for neuron interpretation and manipulation. $\mathsf{NeuronLens}$ maps specific activation ranges within a neuron’s distribution to individual concepts, rather than attributing entire neurons to single concepts. For each concept, $\mathsf{NeuronLens}$calculates a range that covers approximately 98.76% of the activation spectrum, capturing the concept-specific activations while excluding unrelated concepts. Through experiments on encoder-based and decoder-based LLMs across several text classification datasets, we show that $\mathsf{NeuronLens}$. significantly reduces unintended interference by up to 25 percentage points, while maintaining precise control for manipulation of targeted concepts, outperforming existing methods.
In summary, this work makes the following contributions:
  • We uncover that neuronal activations in LLMs form Gaussian-like distributions, with salient neurons exhibiting limited overlap in their activation patterns across concepts.
  • We show that activation ranges within a neuron’s activation spectrum are the true fundamental unit of interpretability, offering a precise framework for neuron-level analysis.
  • We propose $\mathsf{NeuronLens}$, a range-based framework for interpreting and manipulating neuronal activations, which enables fine-grained concept attribution and reduces unintended interference.
@@ 2.1 Preliminaries
Concept. A concept $c\in C$ is a high-level semantic category that groups each input instance (or components of every instance), where $C$ is the set of all concepts. For example, in a language task, a sentence can be categorized into 4 types: declarative, interrogative, imperative, and exclamatory, where each type is a concept. Words of a sentence can also have concepts like noun, verb, adjective, adverb, etc. In this study, we focus on the situation that all input samples are labeled with concepts.
@@
Concept Learning. Given a hidden state vector $\mathbf{h}^{l}$ as input, the associated concept can be the output of an appended neural network (e.g., several fully connected layers). The parameters of this appended neural network can be trained using training samples labeled with concepts.
Table 1: Performance drops relative to Baseline configuration(i.e: unaltered model’s performance) for three techniques: Probeless, Probe, and Max. All values show the difference from Base values. Results are for Emotions dataset on the GPT-2 model. 30% neurons used. The trimmed mean excludes the top and bottom 10% of values.
@@
Class Acc Conf CAcc CConf Acc Conf CAcc CConf Acc Conf CAcc CConf
@@ 2.2 Concept Erasure
To assess the performance of neuronal ranking obtained using attribution methods, concept erasure acts as a critical diagnostic intervention to determine the causal effect of identified neurons for a given concept (Dalvi et al., 2019b; Dai et al., 2022; Dalvi et al., 2019c; Morcos et al., 2018). The core idea is that by manipulating the salient neurons, it is possible to precisely eliminate a specific identified concept while causing minimal disruption to the other concepts learned by the model. This can be formalized as follows: given a concept-learning model $M$ that maps any input instance $x$ (or part of an instance) to a concept $M(x)\in C$, an ideal intervened model $M^{\prime}_{\text{ideal}}$ after erasing a target concept $c\in C$ should satisfy the following property:
$M^{\prime}_{\text{ideal}}(x)\begin{cases}\neq M(x)\quad\text{ if}M(x)=c,\\ =M(x)\quad\text{ if }M(x)\neq c.\end{cases}$
A popular approach of concept erasure in neuronal analysis literature (Dai et al., 2022; Antverg & Belinkov, 2022) is zeroing out specific neurons that are “important” to the target concept.
@@ 2.3 Salient Neurons Extraction
Problem setup and preparation: We use a transformer model to process sentences. Before doing neuron interpretation, we record activations for training samples of different concepts as preparation. Specifically, if we want to interpret neurons of $\mathbf{h}^{l}$ (hidden state vector at layer $l$), we will traverse the training dataset and store the values of $\mathbf{h}^{l}$ and the associated concepts of all samples into a set $H^{l}$. The set $H^{l}$ is further partitioned into $H^{l}_{c}$ for all concepts $c\in C$. Such preparation is commonly used in the literature on neuron interpretation (Dalvi et al., 2019c, b; Antverg & Belinkov, 2022).
Max activations. Frankle & Carbin (2019) extract high neural activations as a saliency ranking metric relying upon the rationale that maximally activating neurons are salient as these neurons play a critical role in controlling the model’s output, highlighting their importance for a concept $c$. Probe analysis. Dalvi et al. (2019b) train a linear classifier on the hidden representations $H^{l}_{c}$ to distinguish between concepts. The learned model weights are then utilized as a saliency ranking. Probeless. Antverg & Belinkov (2022) examine individual neurons, without the need for auxiliary classifiers, using the element-wise difference between mean vectors. Details of these approaches are provided in Appendix C.
Mutual comparison of these approaches is provided in Table 1. From the table, we observed that irrespective of the method used to obtain saliency ranking, a single concept eraser using salient neurons causes deterioration in performance across several concepts. Max activation causes the highest degree of deterioration in the targeted concept while maintaining a comparable deterioration in auxiliary concepts, based on this finding, we adopt max activation ranking for saliency ranking. Moreover, we hypothesize that one reason for such deterioration in overall performance is due to the polysemantic nature of neurons.
Figure 2: Illustrates the percentage overlap of salient neurons identified through maximal activation ranking across concepts (i.e., classes), in all combinations, for the GPT-2 model trained on different datasets. Percentage of salient neurons extracted is 30%.
@@ 3 Polysemanticity
The polysemanticity of neuronal units, salient neurons encode information about multiple concepts, pose a challenge to neural network interpretation and manipulation. In this section, we discuss the degree of polysemanticity in salient neurons in detail.
Figure 3: Neuronal Activation Patterns of six neurons on AG-News dataset. Two on the top are the highest activating neurons, two in the middle are middle-ranked neurons, and two at the bottom are the lowest activating neurons for class 1.
Superposition. Polysemanticity (i.e., superposition of neurons) is primarily thought to arise when a model must represent significantly more features than the available capacity of its representational space or due to the training paradigms used. Limited representation space forces to encode multiple unrelated features within the same neuron to achieve high performance (Anthropic, 2023). Additionally, training paradigms such as subtokenization in tokenization schemes, which aim to reduce vocabulary size to decrease the model size and improve generalization for tokens with the low natural frequency. This results in context-dependent token decomposition, where a single token may be split differently based on its surrounding context (Sennrich et al., 2016). This contextual nature leads to a state of superposition, where multiple semantic meanings are encoded within the same neural activations (Elhage et al., 2022; Meng et al., 2022). Weight initialization can also contribute to polysemanticity, as demonstrated empirically by (Lecomte et al., 2024). Their findings reveal that neurons exhibiting activation patterns equidistant from or similarly close to multiple conceptual representations at initialization can lead to polysemanticity, even when the network’s capacity (i.e., the number of neurons) exceeds the number of concepts to be represented.
Figure 4: Neuronal Activation Patterns comparison of neurons 480 and 675. The plots show class-specific activity patterns with fitted Gaussian curves. Both neurons were part of salient set of all classes when top 5% salient selected on AG-News dataset.
Polysemanticity in salient neurons. Given that salient neurons have a strong causal association with the concept of interest, their tendency to be mono-semantic should be high, but we find that there is a high degree of polysemanticity in salient neurons.
We investigate this by extracting 30% salient neurons (i.e.: Max neurons) for different datasets on the GPT-2 model. The results show that there is a considerable overlap of salient neurons between concepts (classes) as shown in Figure 2. In the case of a two-class dataset IMDB, the overlap of salient neurons, selected by max, is more than 60%. This shows a high degree of polysemanticity.
Consequently, we extrapolate that salient neural representations may exist in a superimposed configuration, wherein a subset of the salient neurons encode information through intricate, overlapping activation patterns. The monolithic attribution paradigm potentially oversimplifies the complex, distributed nature of neuronal activation as can be seen in polysemanticity (Lecomte et al., 2024; Marshall & Kirchner, 2024) where single neuron learns multiple seemingly unrelated concepts and elucidates them at different activation values
@@ 4 Neuronal Activation Patterns
In this section, we analyze the properties of neuronal activations of the salient neurons (including polysemantic) extracted via maximal activation. Our findings indicate that neuronal activations on a concept form a Gaussian-like distribution. We also find that salient neurons have a distinct Gaussian distribution of activations for different concepts with minimal overlap with other concept activations.
Qualitative Evaluation.
@@
To visually demonstrate that neuron activations for a concept $c$ follow a Gaussian-like distribution, we extract model representations as described in Section 2.1. Using saliency ranking $r_{c}$ for a single concept class, we examine neurons from different ranking positions in the GPT-2 model on the AG-News dataset: two top-ranked neurons ($r_{c}\leq 2$), two middle-ranked neurons ($r_{c}\approx d/2$), and two bottom-ranked neurons ($r_{c}\geq d-1$). In Figure 3 and 4 we use Kernel Density Estimation (KDE) to visualize these distributions. Figure 3 reveals that while the activations are Gaussian-like for different concepts, salient neurons demonstrate distinct activation patterns with limited overlap, middle-ranked neurons show a higher degree of overlap than the top ones, whereas non-salient neurons (bottom two) exhibit the highest overlap in their activation distributions. On GPT-2 model.
Additionally in Figure 4, we identify and visualize two distinct types of polysemantic neurons that appear in the salient sets across all classes, when 5% salient set was selected, in the dataset. The first type, exemplified by neuron 480, maintains partially separable activation patterns despite being polysemantic, suggesting some degree of class-specific behavior. In contrast, the second type, represented by neuron 675, exhibits completely overlapping activation patterns across all classes, making it hard to disentangle. To further investigate this phenomenon, Figure 5 presents a broader analysis of neurons from the polysemantic subset, identified using a 5% saliency threshold. By examining these neurons’ behavior across four randomly selected classes (out of 14 total classes), we observe that most polysemantic neurons exhibit a high degree of separability, for some classes – while they respond to multiple classes, they tend to operate in partially separable activation ranges, supporting the possibility of meaningful disentanglement.
Table 2: Statistical analysis of different datasets showing skewness, kurtosis, and Kolmogorov-Smirnov test results.
@@
Quantitative Evaluation. To quantify the effect of Gaussian-like distribution of neurons, for $c\in C$, we perform statistical analysis of activations. We computed the skewness, kurtosis (Joanes & Gill, 1998) and analyzed the normality of neuronal activations using Kolmogorov-Smirnov (KS) test (Massey Jr, 1951). The results for basic distributional properties across all neurons are presented in Table 2. The average skewness is close to 0 across all datasets, indicating strong symmetry (ideal normal distribution: 0), and the average kurtosis is close to 3, nearly identical to the expected value for a normal distribution (3.0), on GPT-2 model.

### v2

@@ Abstract
Interpreting the internal mechanisms of large language models (LLMs) is crucial for improving their trustworthiness and utility. Prior work has primarily focused on mapping individual neurons to discrete semantic concepts. However, such mappings struggle to handle the inherent polysemanticity in LLMs, where individual neurons encode multiple, distinct concepts. Through a comprehensive analysis of both encoder and decoder-based LLMs across diverse datasets, we observe that even highly salient neurons, identified via various attribution techniques for specific semantic concepts, consistently exhibit polysemantic behavior. Importantly, activation magnitudes for fine-grained concepts follow distinct, often Gaussian-like distributions with minimal overlap. This observation motivates a shift from neuron attribution to range-based interpretation. We hypothesize that interpreting and manipulating neuron activation ranges would enable more precise interpretability and targeted interventions in LLMs. To validate our hypothesis, we introduce $\mathsf{NeuronLens}$, a novel range-based interpretation and manipulation framework that provides a finer view of neuron activation distributions to localize concept attribution within a neuron. Extensive empirical evaluations demonstrate that $\mathsf{NeuronLens}$ significantly reduces unintended interference, maintaining precise manipulation of targeted concepts, outperforming neuron attribution.
@@ 1 Introduction
Large language models (LLMs) demonstrate remarkable performance across natural language understanding, generation, and transformation tasks (Brown et al., 2020; Bommasani et al., 2022; Touvron et al., 2023; Raffel et al., 2019). However, the inner workings of LLMs remain largely opaque (Burkart and Huber, 2021), as their representations are distributed across billions of parameters. This lack of interpretability raises critical concerns about reliability, fairness, and trustworthiness, particularly in high-stakes domains such as healthcare, law, and education. To this end, neuron-level interpretability can address these concerns by enabling researchers to uncover how individual neurons encode semantic concepts and contribute to model outputs. With this understanding, researchers can diagnose safety risks (Wei et al., 2024; He et al., 2024), refine model outputs (Meng et al., 2023; Rizwan et al., 2024), optimize efficiency through pruning (Frankle and Carbin, 2019; Haider and Taj, 2021), and steer model’s representations toward desired objectives (Subramani et al., 2022; Li et al., 2023; Rodriguez et al., 2024).
Recent research efforts have made significant progress in neuron-level interpretation by identifying salient neurons that influence model behavior (Dalvi et al., 2019a; Antverg and Belinkov, 2022; Conmy et al., 2023; Marks et al., 2024). Approaches such as maximal activation analysis (Foote et al., 2023; Frankle and Carbin, 2019), Probe-based methods that employ auxiliary classifiers to distinguish between concepts (Dalvi et al., 2019a, b), and the probeless approach bypasses the need for classifiers by directly analyzing neurons (Antverg and Belinkov, 2022). Other techniques include circuit discovery, attributes concepts to groups of interacting neurons (Olah et al., 2020; Conmy et al., 2023), and causal analysis, identify the internal components role in model behavior (Vig et al., 2020).
Figure 1: $\mathsf{NeuronLens}$ leverages distinct, Gaussian-like activation patterns to enable fine-grained concept attribution.
While these methods have advanced our understanding of neurons, they often rely on discrete neuron-to-concept mappings, which assume that entire neurons encode single concepts. However, neurons frequently exhibit polysemanticity; the ability to encode multiple, seemingly unrelated concepts (Lecomte et al., 2024; Marshall and Kirchner, 2024). Given this heterogeneous encoding of concepts, traditional approaches often lead to unintended consequences when manipulating neurons, as changes intended for one concept may inadvertently affect others encoded by the same neuron or suboptimal interpretations of concepts (Sajjad et al., 2022).
@@
This work seeks to address these questions by analyzing the activation patterns of neurons in both encoder-based and decoder-based LLMs. Through statistical and qualitative analysis, we find that neuronal activations for concepts follow Gaussian-like distributions, with distinct patterns for different concepts. Our key insight is that the unit of interpretability lies at a level more fine-grained than the neuron itself. Within a neuron’s activation spectrum, activation ranges corresponding to specific concepts can be used as a finer unit of interpretability. This nuanced perspective enables a more precise approach to neuron interpretation and manipulation, addressing the limitations of traditional, discrete neuron-to-concept mappings.
Building upon these insights, we introduce $\mathsf{NeuronLens}$ visualised in Figure 1, a range-based framework for neuron interpretation and manipulation. $\mathsf{NeuronLens}$ identifies and maps specific activation ranges within a neuron’s distribution to individual concepts, rather than attributing entire neurons to single concepts. For each concept, $\mathsf{NeuronLens}$ calculates a range that covers its activation spectrum, capturing the concept-specific activations while excluding unrelated concepts. Through experiments on encoder-based and decoder-based LLMs across several text classification datasets, we show that $\mathsf{NeuronLens}$ significantly reduces unintended interference by up to 25 percentage points in auxiliary concepts and up to 7x in LLM latent capabilities, while maintaining precise manipulation of targeted concepts, outperforming existing methods.
Our key contributions are: (1) To the best of our knowledge, this is the first work that performs a comprehensive study unfolding polysemantic neurons using activation spectrums. (2) We show that neuronal activations in LLMs form distinct concept level Gaussian-like distributions, with salient neurons exhibiting limited overlap in their activation patterns across concepts. (3) We empirically demonstrate that activation ranges within a neuron’s activation spectrum offer a more precise unit of interpretability, offering a refined framework for neuron-level analysis. (4) We propose $\mathsf{NeuronLens}$, an activation range-based framework for interpreting and manipulating neuronal activations, which enables fine-grained concept attribution and reduces unintended interference compared to neuron level intervention.
@@ 2 Neuron Interpretation Analysis
This section provides an overview of the neuron analysis, methods for extracting salient neurons, and causally validating their saliency.
@@ 2.1 Preliminaries
Concept. A concept $c\in C$ is a high-level semantic category that groups each input instance (or components of every instance), where $C$ is the set of all concepts. For example, in a language task, a sentence can be categorized into 4 types: declarative, interrogative, imperative, and exclamatory, where each type is a concept. Words of a sentence can also have concepts like nouns, verbs, adjectives, adverbs, etc. In this study, we focus on the situation where all input samples are labelled with concepts.
@@
Concept Learning. Given a hidden state vector $\mathbf{h}^{l}$ as input, the associated concept can be the output of an appended neural network (e.g., several fully connected layers). The parameters of this appended neural network can be trained using training samples labelled with concepts.
Metrics. To establish the causal validity of the attribution, we employ two quantitative metrics: prediction accuracy and the model’s predictive probability as a proxy for confidence score. First, baseline measurements of both accuracy and confidence for all concepts $C$ without any intervention (unmodified model) are established. Post-intervention measurements are recorded for the target concept $c$ and auxiliary concepts(other concepts in the dataset) $c^{\prime}$. The effectiveness and precision of attribution are assessed through two key metrics: (1) the magnitude of performance degradation for concept $c$, and (2) the extent of unintended impact on auxiliary concepts $c^{\prime}$. Throughout our analysis, we denote the accuracy and confidence metrics for concept $c$ as Acc and Conf respectively, while corresponding measurements for auxiliary concepts $c^{\prime}$ are represented as CAcc and CConf. For evaluating the effect of the interventions on LLMs latent capabilities, we utilize perplexity (PPL) and MMLU (Hendrycks et al., 2021) zero-shot accuracy.
Table 1: Performance drops relative to Baseline configuration (i.e., unaltered model’s performance) for three techniques: Probeless, Probe, and Max. All values show the difference from Base values. Results are for Emotions dataset on the GPT-2 model using 30% salient neurons of each method. Metrics are detailed in Section 2.1.
@@
 Acc Conf CAcc CConf Acc Conf CAcc CConf Acc Conf CAcc CConf
-0.524 -0.510 -0.086 -0.086 -0.052 -0.036 -0.018 -0.049 -0.735 -0.739 -0.103 -0.103
@@ 2.2 Concept Erasure
To assess the performance of a neuronal attribution, concept erasure acts as a critical diagnostic intervention to determine the causal effect of identified neurons for a given concept (Dalvi et al., 2019b; Dai et al., 2022; Dalvi et al., 2019c; Morcos et al., 2018). The core idea is that if a neuron is salient to a concept, eliminating it should result in the degradation of that concept’s performance while causing minimal disruption to other concepts. This can be formalized as follows: given a concept-learning model $M$ that maps any input instance $x$ (or part of an instance) to a concept $M(x)=c\quad c\in C$, an ideal intervened model $M^{\prime}_{\text{ideal}}$ after erasing a target concept $c\in C$ should satisfy the following property:
$M^{\prime}_{\text{ideal}}(x)=\begin{cases}\neq M(x)&\text{if }M(x)=c,\\ =M(x)&\text{if }M(x)\neq c.\end{cases}$
A popular approach of concept erasure in neuronal analysis literature (Dai et al., 2022; Antverg and Belinkov, 2022) is zeroing out specific neurons that are “important” to the target concept. Other studies have argued that zeroing out neurons is an overly aggressive intervention that can lead to catastrophic degradation in model performance. In Appendix Section D, we provide an ablation comparing different activation interventions for concept erasure.
@@ 2.3 Salient Neurons Extraction
Problem setup and preparation: We record activations for training samples of different concepts to perform neuron interpretation. Specifically, if we want to interpret neurons of $\mathbf{h}^{l}$ (hidden state vector at layer $l$), we traverse the training dataset and store the values of $\mathbf{h}^{l}$ and the associated concepts of all samples into a set $H^{l}$. The set $H^{l}$ is further partitioned into $H^{l}_{c}$ for all concepts $c\in C$. Such preparation is common in the relevant literature (Dalvi et al., 2019c, b; Antverg and Belinkov, 2022).
Max activations. Frankle and Carbin (2019) extract high activations as a saliency ranking metric relying upon the rationale that maximally activating neurons with respect to a concept $c$ are important for that concept. Probe analysis. Dalvi et al. (2019b) train a linear classifier on the hidden representations $H^{l}_{c}$ to predict each concept. The learned model weights are then utilized as a saliency ranking. Probeless. Antverg and Belinkov (2022) examine individual neurons, without the need for auxiliary classifiers, using the element-wise difference between mean vectors. Details of these approaches are provided in Appendix G.
To assess the effectiveness of the attribution methods, we perform neuron masking on the salient neurons identified by each method in a concept erasure task. Table 1 provides the results for this experiment. We observe that irrespective of the method used to obtain saliency ranking, a single concept eraser using salient neurons causes deterioration in performance across several concepts. Max activation causes the highest degree of deterioration in the targeted concept while maintaining a comparable deterioration in auxiliary concepts. Based on this finding, we adopt max activation ranking for saliency ranking. Moreover, we hypothesize that one reason for such deterioration in overall performance is due to the polysemantic nature of neurons.
Figure 2: Overlap of top 30% salient neurons across classes.
@@ 3 Polysemanticity
The polysemanticity of neuronal units, including salient neurons that encode information about multiple concepts, poses a challenge to neural network interpretation and manipulation. In this section, we discuss the degree of polysemanticity in salient neurons in detail.
Polysemanticity often arises when models must represent more features than their capacity allows or due to specific training paradigms. Limited representational space forces neurons to encode multiple unrelated features to maintain performance (Anthropic, 2023). Training methods like subword tokenization, designed to reduce vocabulary size and model complexity, lead to context-dependent token splits, causing neural activations to encode multiple meanings (Sennrich et al., 2016; Elhage et al., 2022; Meng et al., 2022). Additionally, Lecomte et al. (2024) show that even with sufficient capacity, certain weight initializations can induce polysemanticity by placing neurons near multiple conceptual regions.
Polysemanticity in salient neurons. Given that salient neurons have a strong causal association with the concept of interest, their tendency to be mono-semantic should be is high, but we find that there is a high degree of polysemanticity in salient neurons. We investigate this by extracting 30% salient neurons (i.e.: Max neurons) for different datasets on the GPT-2 model. The results show that there is a considerable overlap of salient neurons between concepts (classes) as shown in Figure 2. In the case of a two-class dataset IMDB, the overlap of salient neurons, selected by max, is more than 60%. This shows a high degree of polysemanticity. Consequently, we extrapolate that salient neural representations may exist in a polysemantic configuration, wherein a subset of the salient neurons encode information through intricate activation patterns.
The monolithic attribution paradigm potentially oversimplifies the complex, distributed nature of neuronal activation as can be seen in polysemanticity (Lecomte et al., 2024; Marshall and Kirchner, 2024) where a single neuron learns multiple seemingly unrelated concepts and elucidates them at different activation values.
@@ 4 Neuronal Activation Patterns
In this section, we analyze the properties of neuronal activations of the salient neurons (including polysemantic) extracted via maximal activation. Similar to (Gurnee et al., 2024), our findings indicate that neuronal activations form a Gaussian-like distribution. We further find that salient neurons have a distinct Gaussian distribution of activations for different concepts with limited overlap with other concept activations.
Figure 3: Neuronal Activation Patterns of six neurons on AG-News dataset class 1. Neurons 418 and 447 are the highest activating neurons, neurons 132 and 387 are middle-ranked neurons, and neurons 721 and 365 are the lowest activating neurons.
Figure 4: Comparison of neurons 480 and 675 showing class-specific activation patterns and fitted Gaussian curves. Both neurons were salient across all classes in top 5% on AG-News.
Qualitative Evaluation. To visually demonstrate that neuron activations for a concept $c$ follows a Gaussian-like distribution, we extract model representations as described in Section 2.1. Using saliency ranking $r_{c}$ for a single concept class, we examine neurons from different ranking positions in the GPT-2 model on the AG-News dataset: two top-ranked neurons ($r_{c}\leq 2$), two middle-ranked neurons ($r_{c}\approx d/2$), and two bottom-ranked neurons ($r_{c}\geq d-1$). In Figures 4 and 4, we use Kernel Density Estimation (KDE) to visualize these distributions. Figure 4 reveals that while the activations are Gaussian-like for different concepts, salient neurons demonstrate distinct activation patterns with limited overlap, middle-ranked neurons show a higher degree of overlap than the top ones, whereas non-salient neurons (bottom two) exhibit the highest overlap in their activation distributions.
Additionally in Figure 4, we identify and visualize two distinct types of polysemantic neurons that appear in the salient sets across all classes, when 5% salient set was selected, in the dataset. The first type, exemplified by neuron 480, maintains partially separable activation patterns despite being polysemantic, suggesting some degree of class-specific behavior. In contrast, the second type, represented by neuron 675, exhibits completely overlapping activation patterns across all classes, making it hard to disentangle. To further investigate this phenomenon, Figure 5 presents a broader analysis of neurons from the polysemantic subset, identified using a 5% saliency threshold(top 5% salient neurons selected for a concept). By examining these neurons’ behavior across four randomly selected classes (out of 14 total classes), we observe that most polysemantic neurons exhibit a high degree of separability, for some classes, while they respond to multiple classes, they tend to operate in partially separable activation ranges, supporting the possibility of meaningful disentanglement.
@@
Quantitative Evaluation. To quantify the effect of Gaussian-like distribution of neurons for $c\in C$, we perform statistical analysis of activations. We computed the skewness, kurtosis (Joanes and Gill, 1998) and analyzed the normality of neuronal activations using Kolmogorov-Smirnov (KS) test (Massey Jr, 1951). Table 2 presents the results for distributional properties across all neurons. The average skewness is close to 0 across all datasets, indicating strong symmetry (ideal normal distribution: 0), and the average kurtosis is close to 3, nearly identical to the expected value for a normal distribution (3.0).
Table 2: Skewness, kurtosis, and Kolmogorov-Smirnov test results across various datasets.GPT-2 model